Hi
There is 500g of a 10% salt solution. If water is added to dilute this saline solution to 4% how many grams of water were added.
Thanks

Answer by ankor@dixie-net.com(22740)

There is 500g of a 10% salt solution.
If water is added to dilute this saline solution to 4% how many grams of water were added.
:
let a = amt added to accomplish this
:
the amt of salt does not change, therefore
.04(a + 500) = .10(500)
.04a + 20 = 50
.04a = 50 - 20
.04a = 30
a = 30/.04
a = 750 grams of water needs to added
:
:
Check by finding the amt of salt in each solution, should be equal
.04(750+500) = 50
.10(500) = 50